OpenStudy (anonymous):

A helicopter is at an altitude of 1000 m and is x meters horizontally from a fire. Write the distance, d, from the fire as a function of the horizontal distance, x. I don't need to solve the problem, only write the function

OpenStudy (anonymous):

d^2 = x^2 + (1,000)^2 = x^2 + 1,000,000.